# Centimeter-sized Objects at Micrometer Resolution: Extending Field-of-View in Wavefront Marker X-ray Phase-Contrast Tomography

### v1

@@
Centimeter-sized Objects at Micrometer Resolution: Extending Field-of-View in Wavefront Marker X-ray Phase-Contrast Tomography
††journal: opticajournal††articletype: Research Article‡The authors contributed equally to this work.
[…]
@@
Recent advancements in propagation-based phase-contrast imaging, such as hierarchical imaging, have enabled the visualization of internal structures in large biological specimens and material samples. However, wavefront marker-based techniques, which provide quantitative electron density information, face challenges when imaging larger objects due to stringent beam stability requirements and potential structural changes in objects during longer measurements. Extending the fields-of-view of these methods is crucial for obtaining comparable quantitative results across beamlines and adapting to the smaller beam profiles of fourth-generation synchrotron sources. We introduce a novel technique combining an adapted eigenflat optimization with deformable image registration to address the challenges and enable quantitative high-resolution scans of centimeter-sized objects with micrometre resolution. We demonstrate the potential of the method by obtaining an electron density map of a rat brain sample in diameter using speckle-based imaging, despite the limited horizontal field-of-view of of the beamline (PETRA III, P05, operated by Hereon at DESY). This showcases the ability of the technique to significantly widen the range of application of wavefront marker-based techniques in both biological and materials science research.
1 Introduction
@@
The application of phase-contrast imaging with synchrotron sources has significantly enhanced the exploration of internal structures in materials and biological specimens, offering detailed insights for example into the mechanisms of disease and material structure [1, 2, 3]. A range of different approaches have been developed and successfully applied to extract the phase information imparted onto the X-ray wavefront by the sample, including e.g. propagation-based imaging[4, 5], ptychographic methods [6], a Fokker-Planck based approach[7, 8], and speckle-based imaging [9]. Currently, methods for imaging large objects at the highest possible resolution are an active area of research, because they can provide a better understanding of biological processes and material properties [10]. They are becoming even more relevant with the advent of fourth generation synchrotron sources, which typically feature very small angular source widths [11], leading to smaller fields of view at the same measurement distance. Recently, hierarchical propagation-based imaging has been used successfully to obtain morphological information of different human organs down to the micrometre level [12]. For this purpose, a large number of individual scans is acquired and stitched together to yield information on the complete organ.
@@
A similar approach is currently difficult to replicate for wavefront-marker-based techniques, but would allow retrieving quantitative information on the electron density of the large object while also not requiring the single-material assumption (i.e. a fixed -ratio). In these techniques, a wavefront marker is placed in the beam path to create a reference pattern, which is then distorted by the object. Processing approaches such as cross correlation [13] or least-squares minimization (e.g. Unified Modulated Pattern Analysis (UMPA) [14]) then allow retrieving three different signals—attenuation, phase-contrast and dark-field—from multiple tomograms taken at different marker positions. For the purpose of accurate signal retrieval, the beam profile needs to remain constant for both the reference and the sample scan, i.e. the source must remain stable over the time horizon of the measurement—a condition assumed in conventional flat-field correction [15]. For most synchrotron radiation sources, however, the beam profile varies significantly over time (e.g. due to beamline component motion or the electron beam orbit control). These discrepancies between the reference image and the conditions during the measurement may cause the wavefront marker (e.g. a grid pattern [16]) to remain visible in the retrieved projections, compromising the image quality and quantitative accuracy of the imaging method.
@@
To address this problem, approaches like a sample/diffusor drift correction for the case of UMPA [17] and an eigenflat approach have been developed. In the latter approach, a group of acquired flat-field images is decomposed into eigencomponents. A combination of these eigencomponents is then fitted to each projection separately to provide a synthetic flat-field image [18]. This method has also been successfully applied to phase-contrast scans using speckle-based imaging, but currently requires the presence of a sample-free region within the scan to accurately capture the displacement of the wavefront markers [19]. When acquiring multiple tomograms at different positions of the same object, only the tomograms capturing the edges of the object include such a sample-free region. For the inner tomograms, the entire field-of-view is covered by the object and the current eigenflat method is therefore not applicable.
@@
An additional challenge for stitching multiple such tomograms are possible deformations experienced by the object due the prolonged exposure to intense radiation. Such deformations may not be accurately described by assuming rigid transformations (translation and rotation), as is typically the case for conventional stitching techniques used for panorama creation [20] or medical diagnosis [21]. For this problem, we therefore propose the use of deformable image registration, a technique applied successfully for example in image-guided surgery [22] and radiation therapy [23].
@@
In response to both of the challenges faced by wavefront-marker-based techniques, we present a novel approach combining an eigenflat method developed for this scenario with deformable image registration. In this way, we allow for an arbitrary expansion of the total field-of-view by combining multiple tomograms acquired at different positions to cover the entire sample. This means it becomes possible to obtain electron density maps for objects several centimeters in size with micrometer resolution without strict requirements on the stability of the beam profile or sample over time. Due to the quantitativeness of the method, the results are better comparable over time and at different beamlines.
@@
As an example application, the method is used to retrieve an electron density map of the full horizontal extent of a rat brain in diameter at the Hereon-operated beamline P05 at PETRA III (DESY, Hamburg), despite the limited horizontal FOV. This scan uses Talbot Array Illuminators as wavefront markers, though the method is also applied towards scans using sandpaper in the supplementary document.
2 Methodology
2.1 Scan Acquisition
The general scanning strategy for samples that do not fit into a single FOV is illustrated in Figure 1a. The approach is applicable to any number of recorded subregions, but is described for the case of scans as visualized by Figure 1b in the following.
@@ 2.1 Scan Acquisition
At first, the left outer, inner, and right outer regions of the sample are scanned separately, resulting in three sets of computed tomography (CT) projections. The iso-centers of these scans are set fixed and the difference between the scans is equivalent to a change in the horizontal offset of the detector and source. The horizontal offsets of the detector and source are chosen so that there is a small overlap between projection images from different sample parts, indicated as left and right overlap respectively in Figure 1b.
2.2 Processing
@@
The initial processing is shown in Figure 2a. The outer projection images are acquired such that there are sample-free areas at the borders. Then the eigenflat-field correction as first described by [18] is applied: Given a set of recorded flat-fields with pixels each, the mean of all flat-fields is first subtracted from each to obtain the centered flat-field matrix :
@@
The eigenflat-fields then refer to the eigenvectors of the covariance matrix . For each image, a matching flat-field is calculated by least-squares optimization of the difference between pixels in a sample-free region of the image and the corresponding region in the flat-field , where are the weights subject to optimization:
@@
The optimized flats are used in the speckle-matching algorithm to obtain the differential phase images in and direction. In a first step, the inner projection images of the sample are also input into the speckle-matching algorithm together with an averaged reference image created by computing the average of all originally recorded flat-field images.
@@
The data is then processed using an implementation of UMPA [24], yielding differential phase images in two orthogonal directions for all three parts of the image. The inner projections contain grid-shaped artifacts due to the time-varying nature of the beam, which does not allow to remove the wavefront marker pattern completely using an averaged flat approach. Before an adapted eigenflat correction may be applied to resolve this problem, the three image parts first need to be registered precisely.
2.3 Registration
@@
Structural changes may be experienced by the samples due to the applied radiation dose. These include thermal expansion or changes in chemical composition due to radiolysis. This, and different beam profiles at different positions within the FOV, mean that the pair of images to be registered may not be perfectly aligned if the images are merely rigidly shifted based on the horizontal offsets. To cope with the potentially complicated deformation that may occur during the course of the scans, a deformable image registration algorithm (ImFusion GmbH, Munich, Germany [25]) is applied. A free-form deformation (FFD) model [26] with B-Splines interpolation is used. The deformation model is modified by a control point grid. For each pixel inside the grid, its displacement is determined by the displacements of the nearest 4 control points. The resulting deformation field is continuous. Aiming at finding the optimal deformation model, the deformable registration is equivalent to an optimization problem. During the optimization, the displacements of the control points are the variables to be optimized, and the similarity metric between the images is to be maximized.
@@
The deformable registration is demonstrated in Figure 2b. The inner image is assumed to be stationary and left undeformed, while the outer regions receive deformation models. This particular choice creates a computational advantage for the adapted eigenflat-approach introduced in the next subsection. The deformable image registration is then performed on either the differential phase images in the - or the -direction for all projections. We perform the registration consecutively for all the projections. To encourage smoothness between adjacent projections, as well as improve computational efficiency, we use the optimal control point displacements from the projection at the current angle as the initial guess of the projection at the next angle.
2.4 Adapted Dynamic Flat-field Correction
@@
In a next step, eigenflats for the inner scan are calculated. However, the matching flat-field images cannot be obtained directly via the approach described in the previous section: The specific combination of weights (see Eq. (2)) cannot be computed because no sample-free region to be used for matching is present. A different optimization problem targeting the optimal weights for the eigenflats of the inner projection is therefore set up, as shown in Figure 2c.
@@
for conciseness. We can then use the pixel values and from the inner and right outer scans respectively to create a difference image . Inside a region , which we define as the overlap region(s) between the inner and outer scans, is approximated as follows
where we have decomposed the difference between the two images into a residual speckle pattern and a pixel-dependent noise term . In an ideal case in which a perfectly matching flat-field has been used for the correction, the term should equal 0, since the wavefront marker is successfully removed by the speckle-tracking algorithm. Neglecting the noise term, we may therefore use the homogeneity of the difference image as a measurement of the flat-field correction quality. We choose the total variation (TV) for this purpose, since it is a well established metric for evaluating the spatial variation and the smoothness of an image [27]. For an image it is defined as:
where and denote the pixel indices in horizontal and vertical direction of the 2D image.
The TV of the difference image is now computed and returned as the value of the cost function of the optimization. The cost function is defined as:
which is eventually minimized to find the optimal weights , as mentioned in Eq. 2.
Using the determined weights for the eigenflats, new flat-field images are generated and used to retrieve the final, corrected differential phase images of the middle scan. The left outer, inner and right outer differential phase-contrast images are then transformed and deformed using the information of the horizontal offsets and the deformation model previously determined by the deformable registration.
2.5 Stitching
Finally, the overlapping regions of the images are blended using a distance weighted sum
@@ 2.5 Stitching
Here, the intensity of a pixel in the overlapping region is computed from the pixel intensity of the inner image and the right outer image . The weight depends on width of the overlapping region and the distance of the current pixel to the left boundary of the overlapping region .
The three projections are stitched using the same method to form one projection with enlarged FOV. The stitched differential phase images in the and direction are then integrated using a 2D Fourier integration approach [28] to retrieve the final integrated phase signal, which is used for the three-dimensional reconstruction, yielding an electron density map.
2.6 Cascading Approach for Arbitrary Number of Projections
It is to be noted that the proposed method is applicable to an arbitrary number of projections instead of three as shown in Figure 1b. The correction will then be performed consecutively from outer to inner projections. For a projection , where , the overlapping region with previously corrected projection is considered. The cost function for the optimization on projection is defined as
@@ 2.6 Cascading Approach for Arbitrary Number of Projections
3 Experiments
3.1 Beamline Parameters
@@
The imaging beamline P05 uses X-rays from an undulator located d = from the microtomography endstation. A double-crystal silicon monochromator provides a highly monochromatic beam () with an energy of . The beam is approximately parallel ( at 10 keV [29]). A CMOS camera developed in collaboration with the Karlsruhe Institute of Technology (KIT) based on the CMOSIS CMV 20000 sensor is positioned behind the sample. It has an array of pixels with a physical pixel size of , which leads to an effective pixel size of using a 5x magnification lens. The objective is coupled to a scintillator.
@@
The wavefront marker is a 2D phase shifting grating called a Talbot Array Illuminator [16] positioned in front of the sample. The grating has a period of and produces a phase shift of using a hexagonal grating structure. The duty cycle is and specifies the ratio of hole spacing to hole radius. The gratings were fabricated on thick silicon wafers using deep reactive ion etching to produce round holes of suitable depth. At fractional Talbot distances of , these TAIs exhibit a focusing effect with a theoretical compression ratio of 1:3 in each direction, allowing high visibility to be achieved. The gratings are stepped using a 2D piezoelectric stepper.
3.2 Sample
@@
The sample investigated during this work is a whole rat brain. The animal housing and organ removal was carried out at Helmholtz-Zentrum Munich following the European Union guidelines 2010/63. The procedure was conducted in compliance with the ethical standards of the institution and approved by the responsible governmental body. Specifically, the experiments were performed under the license number ROB-55.2-2532.Vet_02-21-133. After removal, the sample was fixated in Formaldehyde. Then, the sample was dehydrated using an ethanol series starting with 50%(vol/vol) ethanol per water up to 100%(vol/vol) ethanol in steps of 10% for one hour each. Afterwards, the sample was stained in an eosin stain on ethanol basis, with 0.6%(wt/vol) eosin on 100% ethanol. The ethanol-based staining was used due to the high lipid ratio in the brain tissue. The brain contains a MENX-associated mutation [30].
3.3 Measurement Protocol

### v2

# Extending the field of view in modulation-based X-ray phase microtomography

@@
Abstract
@@
Recent advances in propagation-based phase-contrast imaging, such as hierarchical imaging, have enabled the visualization of internal structures in large biological specimens and material samples. However, modulation-based techniques, which provide quantitative electron density information, face challenges when imaging larger objects due to stringent beam stability requirements and detector distortions. Extending the field of view of these methods is crucial for obtaining comparable quantitative results across beamlines and adapting to the smaller beam profiles of fourth-generation synchrotron sources. We introduce a novel image processing technique combining an eigenflat optimization with deformable image registration to address the challenges and enable quantitative high-resolution scans of centimeter-sized objects with multiple-micrometer resolution. We demonstrate the potential of the method by obtaining an electron density map of a rat brain sample 15 mm in diameter despite the limited horizontal field of view of 6 mm of the beamline. This showcases the technique’s ability to significantly widen the range of applications of modulation-based techniques in both biological and materials science research.
I Introduction
@@
The application of phase-contrast imaging at synchrotron sources has significantly enhanced the exploration of internal structures in materials and biological specimens, offering detailed insights, for example, into the mechanisms of disease and material structure [42, 20, 41]. A range of different approaches has been developed and successfully applied to extract the phase information imparted onto the X-ray wavefront by the sample, including e.g. propagation-based imaging[23, 28], ptychographic methods [25], Fokker-Planck based approaches [22, 18], and speckle-based imaging [49] – the latter being more generally referred to as modulation-based imaging to include structured diffusors [27, 36]. Methods for imaging large objects at the highest possible resolution are currently an active area of research, because they can provide a better understanding of biological processes and material properties [11, 32, 2]. They are becoming even more relevant with the advent of fourth-generation synchrotron sources, which typically feature very small angular source widths [9], leading to smaller fields of view at the same measurement distance. Recently, hierarchical propagation-based imaging has successfully been used to obtain morphological information of different human organs down to the micrometer level [46]. For this purpose, projections are acquired at multiple positions and concatenated before reconstruction to yield information on the complete organ.
@@
A similar approach is currently difficult to replicate for modulation-based techniques, but would allow retrieving quantitative information on the electron density of the large object while also not requiring the single-material assumption (i.e., a fixed -ratio). In these techniques, a modulator is placed in the beam path to create a reference pattern, which the object then distorts. Processing approaches such as cross-correlation [4] or least-squares minimization (e.g. Unified Modulated Pattern Analysis (UMPA) [48]) then allow retrieving three different signals – attenuation, phase-contrast and dark field – from multiple tomograms taken at different marker positions. For the purpose of accurate signal retrieval, the beam profile needs to remain constant for both the reference and the sample scan, i.e. the source must remain stable over the time horizon of the measurement – a condition assumed in conventional flat-field correction [40]. For most synchrotron radiation sources, the beam profile varies significantly over time (e.g. due to beamline component motion or the electron beam orbit control). These discrepancies between the reference image and the conditions during the measurement may cause the modulator (e.g., a grid pattern [13]) to remain visible in the retrieved projections, compromising the image quality and quantitative accuracy of the imaging method.
@@
To address this problem, approaches like a sample/diffusor drift correction for the case of UMPA [34, 35] and an eigenflat approach have been developed. In the latter approach, a group of acquired flat-field images is decomposed into eigencomponents. A combination of these eigencomponents is then fitted to each projection separately to provide a synthetic flat-field image [21]. This method has also been successfully applied to phase-contrast scans using modulation-based imaging, but currently requires the presence of a sample-free region within the scan to accurately capture the displacement of the modulator [29]. When acquiring multiple tomograms at different positions of the same object, only the tomograms capturing the edges of the object include such a sample-free region. For the inner tomograms, the entire field of view is covered by the object and the current eigenflat method is therefore not applicable.
@@
An additional challenge for stitching multiple such tomograms are possible distortions of the projection images due to the time-varying beam profile and the spatially varying detector distortion. Such distortions may not be accurately described by assuming rigid transformations (translation and rotation), as is typically the case for conventional stitching techniques used for panorama creation [14] or medical diagnosis [1]. For this problem, we therefore propose the use of deformable image registration, a technique applied successfully for example in image-guided surgery [31] and radiation therapy [7].
@@
In response to the mentioned challenges faced by modulation-based techniques, we present a novel approach combining an eigenflat method developed for this scenario with deformable image registration. In this way, we allow for an arbitrary expansion of the total field of view by combining multiple tomograms acquired at different positions to cover the entire sample. This means it becomes possible to obtain electron density maps for objects several centimeters in size with multiple-micrometer resolution without strict requirements on the stability of the beam profile or sample over time. Due to the quantitative nature of the method, the results are more comparable over time and at different beamlines.
@@
As an example application, the method is used to retrieve an electron density map of the full horizontal extent of a rat brain in diameter at the beamline P05, operated by Hereon, at PETRA III (DESY, Hamburg), despite the limited horizontal FOV. This scan utilizes Talbot array illuminators as modulators; however, we also apply the method to data acquired with a sandpaper modulator, as described in the appendix.
II Methodology
II-A Scan acquisition
@@
The general scanning strategy for samples that do not fit into a single FOV is illustrated in Figure 1(a). The approach is applicable to any number of recorded subregions, but is described for the case of scans as visualized in Figure 1(b) in the following. At first, full tomographic data sets of the left outer, inner, and right outer regions of the sample are acquired separately and in this order, resulting in three sets of computed tomography (CT) projections. The iso-centers of these scans are set fixed and the difference between the scans is equivalent to a change in the horizontal offset of the detector and source. The horizontal offsets of the detector and source are chosen such that there is a small overlap between projection images from different sample parts, indicated as left and right overlap respectively in Figure 1(b). The overlap size should be chosen on the order of of the single field of view.
II-B Processing
@@
The initial processing is shown in Figure 2(a). The outer projection images are acquired such that there are sample-free areas at the borders. The eigenflat-field correction, as first described in Ref. [21], is then applied: Given a set of recorded flat-field images with pixels each, the mean of all flatfields is first subtracted from each to obtain the centered flatfield matrix :
@@
The eigenflat-fields are defined as the eigenvectors of the covariance matrix . For each image, a matching flatfield is calculated by least-squares optimization of the difference between pixels in a sample-free region of the image and the corresponding region in the flatfield , where are the weights subject to optimization:
@@
The optimized flats are used in the modulation-matching algorithm to obtain the differential phase images in - and -direction. In a first step, the inner projection images of the sample are also input into the modulation-matching algorithm together with a mean reference image created by computing the average of all originally recorded flat-field images.
@@
The data are then processed using an implementation of UMPA [10], yielding differential phase images in two orthogonal directions for all three parts of the image. The inner projections contain grid-shaped artifacts due to the time-varying nature of the beam, which does not allow to remove the modulator pattern completely using a mean flat approach. Before an adapted eigenflat correction can be applied to resolve this problem, the three image parts first must be registered precisely.
II-C Registration
@@
The projection images may exhibit structural changes due to detector distortion at different positions. These effects, together with different beam profiles at various positions within the FOV, mean that the pair of images to be registered may not be perfectly aligned if the images are merely rigidly shifted based on the horizontal offsets. To cope with these potentially complicated deformations, we apply a deformable image registration algorithm (ImFusion GmbH, Munich, Germany [15]). A free-form deformation (FFD) model [33] with B-Splines interpolation is used. The deformation model is modified by a control point grid. For each pixel inside the grid, its displacement is determined by the displacements of the 4 nearest control points. The resulting deformation field is continuous. Aiming to find the optimal deformation model, the deformable registration is equivalent to an optimization problem. During the optimization, the displacements of the control points are the variables to be optimized, and the similarity metric between the images is to be maximized.
@@
The deformable registration is demonstrated in Figure 2(b). The inner image is assumed to be stationary and left undeformed, while the outer regions receive deformation models. This particular choice creates a computational advantage for the adapted eigenflat-approach introduced in the next subsection. The deformable image registration is then performed on either the differential phase images in the - or the -direction for all projections. We perform the registration consecutively for all the projections. To encourage smoothness between adjacent projections, as well as improve computational efficiency, we use the optimal control point displacements from the projection at the current angle as the initial guess of the projection at the next angle.
II-D Dynamic flat-field correction
@@
In the next step, eigenflats for the inner scan are calculated. However, the matching flat-field images cannot be obtained directly via the approach described in the previous section: The optimal weights (see Eq. (2)) cannot be computed directly because no sample-free region for matching is present. A different optimization problem targeting the optimal weights for the eigenflats of the inner projection is therefore set up, as shown in Figure 2(c).
@@
for conciseness. We can then use the pixel values and from the inner and right outer scans respectively to create a difference image . We note here that refers to the outer image after application of the deformable registration. Inside a region , defined as the overlap region(s) between the inner and outer scans, is approximated as follows:
Here, we have decomposed the difference between the two images into a residual modulation pattern and a pixel-dependent noise term . This decomposition is valid provided that a sufficiently good match has been achieved between the inner and outer image regions during registration, such that any remaining difference arises from the modulator pattern or noise rather than displaced image features. In an ideal case in which a perfectly matching flat-field has been used for the correction, the term should equal 0, since the modulator pattern is successfully removed by the modulation-tracking algorithm. Neglecting the noise term, we may therefore use the homogeneity of the difference image as a measurement of the flat-field correction quality. We choose the total variation (TV) for this purpose, since it is a well-established metric for evaluating the spatial variation and the smoothness of an image [8]. For a two-dimensional image , it is defined as
where and denote the pixel indices in the horizontal and vertical direction of the image. The TV of the difference image is then computed and returned as the value of the cost function of the optimization. The cost function is defined as
which is eventually minimized to find the optimal weights , as mentioned in Eq. (2).
Using the determined weights for the eigenflats, new flat-field images are generated and used to retrieve the final, corrected differential phase images of the middle scan. The left outer, inner, and right outer differential phase-contrast images are then transformed and deformed using the information from the horizontal offsets and the deformation model previously determined by the deformable registration.
II-E Stitching
Finally, we blend the overlapping regions of the images using a distance-weighted sum
@@ II-E Stitching
Here, the intensity of a pixel in the overlapping region is computed from the pixel intensity of the inner image and the right outer image . The weight depends on the width of the overlapping region and the distance of the current pixel to the left boundary of the overlapping region .
The three projections are stitched using the same method to form one projection with an enlarged FOV. The stitched differential phase images in the - and -direction are then integrated using a 2D Fourier integration approach [17] to retrieve the final integrated phase signal. This signal is then reconstructed into a three-dimensional electron density map using a Feldkamp-Davis-Kress (FDK) algorithm, as implemented in the Xaid software (Mitos GmbH, Garching, Germany).
II-F Cascading approach for arbitrary number of projections
We note here that the proposed method is applicable to an arbitrary number of projections instead of three, as shown in Figure 1(b). The correction will then be performed consecutively from outer to inner projections. For a projection , where , the overlapping region with the previously corrected projection is considered. The cost function for the optimization on projection is defined as
@@ II-F Cascading approach for arbitrary number of projections
III Experiments
III-A Beamline parameters
@@
The imaging beamline P05 uses X-rays from an undulator located from the microtomography endstation. A double-crystal silicon monochromator provides a highly monochromatic beam () with an energy of . The beam is approximately parallel with a beam divergence of at [47]. A CMOS camera based on the CMOSIS CMV 20000 sensor [19], developed in collaboration with the Karlsruhe Institute of Technology (KIT), is positioned behind the sample. It features an array of pixels with a physical pixel size of , which leads to an effective pixel size of using a 5-fold magnification lens. The objective is coupled to a -thick scintillator.
@@
The modulator is a two-dimensional phase-shifting grating referred to as a Talbot array illuminator [13]. It is positioned in front of the sample. The modulator has a period of and introduces a phase shift of to parts of the beam in a hexagonal lattice structure. The duty cycle, which specifies the ratio of hole spacing to hole radius, is . The modulator was fabricated on -thick silicon wafers using deep reactive ion etching to produce round holes of suitable depth. At fractional Talbot distances of , the modulators exhibit a focusing effect with a theoretical compression ratio of 1:3 in each direction, allowing high visibility to be achieved. The grating is stepped using a piezoelectric stepper capable of movement in the two directions orthogonal to the beam path.
III-B Sample
@@
The sample investigated during this work is a whole rat brain. Animal housing and organ removal were carried out at Helmholtz-Zentrum Munich in accordance with the European Union guidelines 2010/63. The procedure was conducted in compliance with the ethical standards of the institution and approved by the responsible governmental body. Specifically, the experiments were performed under the license number ROB-55.2-2532.Vet_02-21-133. After removal, the sample was fixed in formaldehyde. The sample was then dehydrated using an ethanol series starting with (vol/vol) ethanol per water up to (vol/vol) ethanol in steps of for one hour each. Afterwards, the sample was stained with an eosin stain on an ethanol basis, using (wt/vol) eosin in ethanol. The ethanol-based staining was used due to the high lipid ratio in the brain tissue. The brain contains a MENX-associated mutation [24].
III-C Measurement protocol